OpenStudy (anonymous):

percent of increase 120 to 200

OpenStudy (anonymous):

40

hero (hero):

You're still trying to figure these out @AngelDreamGurl ?

OpenStudy (anonymous):

YES *puppy dog face* LOL

hero (hero):

What seems to be the trouble with following the formula I gave you?

OpenStudy (anonymous):

its kind of confusing

hero (hero):

What exactly is so confusing about it? All you do is plug in the numbers

OpenStudy (anonymous):

what is the formula, Hero? Wish I had one of those when I was in school

hero (hero):

\[\% \Delta = \frac{|a - b|}{a} \times 100\] a = old number b = new number

OpenStudy (anonymous):

on this one, I would say to myself, self, 80 is what percent of 120?

hero (hero):

applying absolute value symbols are important for this one

OpenStudy (anonymous):

yes, and those are some daunting symbols if you aren't in precalc or beyond

hero (hero):

\[\% \Delta = \frac{|120 - 200|}{120} \times 100\]

OpenStudy (anonymous):

66.6

OpenStudy (anonymous):

OHHHHHHHHHHHHHHHHHHHHHH i missed the absolute value sign

hero (hero):

\[\% \Delta \space \text{= percent change}\]

hero (hero):

There's nothing daunting about that

OpenStudy (anonymous):

delta and the absolute value signs are... at first

hero (hero):

@tutor_guy , did you want to say something? Probably to dispute my formula

OpenStudy (anonymous):

No, I actually gave @AngelDreamGurl similar formula as yours. :)))

OpenStudy (anonymous):

that's the formula all right, been using it over 50 years now, just never saw it written formally like that before. I just knew that increase was change/start

OpenStudy (anonymous):

and that percent always mean to divide by 100